xx TestSchedule

TestSchedule

Tom Goodale goodale@cct.lsu.edu

Date

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters 7 Interfaces

General Implements: testschedule

8 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestSchedule. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x testschedulexcctkxstartup x test to see if function is placed in schedule and run at cctkxstartup x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxRECOVERxPARAMETERS x testschedulexcctkxrecoverxparameters x test to see if function is placed in schedule and run at cctkxrecoverxparameters CCTKxCHECKPOINT x testschedulexcctkxcheckpoint x test to see if function is placed in schedule and run at cctkxcheckpoint x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPRESTEP x testschedulexcctkxprestep x test to see if function is placed in schedule and run at cctkxprestep x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxEVOL x testschedulexcctkxevol x test to see if function is placed in schedule and run at cctkxevol x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTSTEP x testschedulexcctkxpoststep x test to see if function is placed in schedule and run at cctkxpoststep x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTRESTRICT x testschedulexcctkxpostrestrict x test to see if function is placed in schedule and run at cctkxpostrestrict x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTREGRID x testschedulexcctkxpostregrid x test to see if function is placed in schedule and run at cctkxpostregrid CCTKxANALYSIS x testschedulexcctkxanalysis x test to see if function is placed in schedule and run at cctkxanalysis x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxTERMINATE x testschedulexcctkxterminate x test to see if function is placed in schedule and run at cctkxterminate x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxSHUTDOWN x testschedulexcctkxshutdown x test to see if function is placed in schedule and run at cctkxshutdown x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxWRAGH x testschedulexcctkxwragh x test to see if function is placed in schedule and run at cctkxwragh x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPARAMCHECK x testschedulexcctkxparamcheck x test to see if function is placed in schedule and run at cctkxparamcheck x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxBASEGRID x testschedulexcctkxbasegrid x test to see if function is placed in schedule and run at cctkxbasegrid CCTKxINITIAL x testschedulexcctkxinitial x test to see if function is placed in schedule and run at cctkxinitial x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTINITIAL x testschedulexcctkxpostinitial x test to see if function is placed in schedule and run at cctkxpostinitial x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxRECOVERxVARIABLES x testschedulexcctkxrecoverxvariables x test to see if function is placed in schedule and run at cctkxrecoverxvariables x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxPOSTxRECOVERxVARIABLES x testschedulexcctkxpostxrecoverxvariables x test to see if function is placed in schedule and run at cctkxpostxrecoverxvariables x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxCPINITIAL x testschedulexcctkxcpinitial x test to see if function is placed in schedule and run at cctkxcpinitial x

11 x 11 Language: 11 c
11 x 11 Type:  11 function